# PRISM: Leveraging Prototype Patient Representations with Feature-Missing-Aware Calibration for EHR Data Sparsity Mitigation

### v5

@@ Abstract
Electronic Health Records (EHRs) contain a wealth of patient data; however, the sparsity of EHRs data often presents significant challenges for predictive modeling. Conventional imputation methods inadequately distinguish between real and imputed data, leading to potential inaccuracies of patient representations. To address these issues, we introduce PRISM, a framework that indirectly imputes data by leveraging prototype representations of similar patients, thus ensuring compact representations that preserve patient information. PRISM also includes a feature confidence learner module, which evaluates the reliability of each feature considering missing statuses. Additionally, PRISM introduces a new patient similarity metric that accounts for feature confidence, avoiding overreliance on imprecise imputed values. Our extensive experiments on the MIMIC-III, MIMIC-IV, PhysioNet Challenge 2012, eICU datasets demonstrate PRISM’s superior performance in predicting in-hospital mortality and 30-day readmission tasks, showcasing its effectiveness in handling EHR data sparsity. For the sake of reproducibility and further research, we have made the code publicly available at https://github.com/yhzhu99/PRISM.
footnotetext: Corresponding author.
1 Introduction
@@
Electronic Health Records (EHR) have become indispensable in modern healthcare, offering a rich source of data that chronicles a patient’s medical history. Over recent years, machine learning techniques have gained significant attention for their ability to leverage time-series EHR data, which represented as temporal sequences of high-dimensional clinical variables [1], can significantly inform and enhance clinical decision-making. Such applications range from predicting the survival risk of patients [2, 3] to forecasting early mortality outcomes [4, 5, 6, 7].
@@
Working with time-series EHR data presents challenges due to its inherent sparsity. Factors such as data corruption [8], expensive examinations [9], and safety considerations [10] result in missing observations; for instance, not all indicators are captured during every patient visit [11]. Imputed values, while necessary, are not genuine reflections of a patient’s condition and can introduce noise, diminishing model accuracy [12]. Given that most machine learning models cannot process NaN (Not a Number) inputs, this sparsity necessitates imputation, complicating EHR predictive modeling. While most existing works tackling EHR sparsity have tried to perform the imputation task directly on raw data, based on modeling the health status trajectory of the whole training set, this approach is similar to that of matrix completion methods, such as MICE [13], non-negative matrix factorization (NMF) [14], and compressed sensing [15]. However, they fail to capture the temporal interactions of longitudinal EHR data for each patient. Also, previous EHR-specific models with strategies of recalibrating patient representations based on attentive feature importance [16, 17] exhibit an issue of inadvertently prioritizing imputed features, potentially introducing inaccuracies [18]. Thus, addressing the sparsity in time-series EHR data requires a focus on capturing relevant feature representations across visits within a patient or among patients that are essential for predictive purposes. The primary objective is to discern and highlight crucial features while diminishing the impact of irrelevant, redundant, or missing ones.
@@
Intuitively, incorporating knowledge from similar patients as an indirect method of imputation has the potential to enhance patient representations in the context of sparse EHR data. Such a knowledge-driven approach mirrors the real-world clinical reasoning processes, harnessing patterns observed in related patient cases [19]. Regarding the identifying similar patient process, however, existing models also face a significant limitation: they cannot distinguish between actual and imputed data [3, 12]. Consider two patients who have the same lab test feature value. For patient A, this value originates from the actual data, while for patient B, it is an imputed value. Current similarity metrics, whether they are L1, L2 distance, cosine similarity [20], handcrafted metrics [21], or learning-based methods [19], interpret these values identically. This results in a potentially misleading perception of similarity.
@@
Given these insights, we are confronted with a pressing challenge: How can we effectively mitigate the sparsity issue in EHR data caused by missing recorded features while ensuring a compact patient representation that preserves patient information? [22]
@@
  • Methodologically, we propose PRISM, a framework for learning prototype representations of similar patients, designed to mitigate EHR data sparsity. We design the feature confidence learner that evaluates and calibrates the reliability of each feature by examining its absence and associated confidence level. We also introduce the confidence-aware prototype patient learner with enhanced patient similarity measures that differentiates between varying feature confidence levels. Compared to existing SOTA baselines, i.e. GRASP [3] and M3Care [12], PRISM provides a refined feature calibration, and further missing-aware similarity measure helps to identify more related patient representation, with missing feature status elaborately taking into account.
@@ 2 Related Work
2.1 Direct Imputation
Direct imputation methods aim to estimate missing features or incorporate missing information directly. Traditional matrix imputation techniques, such as MICE [13], non-negative matrix factorization (NMF) [14], compressed sensing [15], or naive zero, mean, or median imputation, rely on similar rows or columns to fill in missing data. However, these methods often operate under the assumption that patient visits are independent and features are missing at random. GRU-D [11] adopts a more targeted approach by introducing missing statuses into the GRU network. By utilizing time interval and missing mask information, GRU-D treats missing data as “Informative Missing”. Extending these capabilities, ConCare [2] and AICare [17] first apply directly imputed EHR data, then incorporate multi-head self-attention mechanisms to refine feature embeddings. This ensures contextual relevance across diverse healthcare situations, regardless of data completeness.
2.2 Indirect Imputation
As demonstrated in GRASP [3] and M3Care [12], they emphasize the use of similar patient representations to derive meaningful information with the insight that the information observed from similar patients can be utilized as guidance for the current patient’s prognosis [3]. However, accurately measuring patient similarity is intrinsically challenging, especially when features might be imputed with potentially misleading information. Many traditional works, such as  [20] and  [21], have resorted to fixed formulas like cosine similarity and Euclidean distance to gauge patient similarity. While these methods are straightforward, they often suffer from scalability and performance limitations. A more dynamic approach is seen in  [19], which adopts metric learning with triplet loss. This technique focuses on learning the relative distances between patients, where distances have an inverse correlation with similarity scores.
@@ 2.2 Indirect Imputation
3 Problem Formulation
3.1 EHR Datasets Formulation
@@
3.2 Predictive Objective Formulation
@@ 3.3 Notation Table
Table 1: Notations symbols and their descriptions
@@
4 Methodology
4.1 Overview
Figure 1: Overall model architecture of our proposed method PRISM. “Rep.” means “Representation”.
@@ 4.1 Overview
4.2 Feature-Isolated Embedding Module
@@
4.2.1 Static Features Embedding
@@
$\begin{array}[]{cc}\bm{h_{s_{i}}}=\mathrm{MLP}_{i}(\bm{s_{i}}),&i=1,2,\cdots,S\end{array}$ (1)
@@
4.2.2 Dynamic Features Embedding
To ensure that each feature’s individual statistics, e.g. missing status can be incorporated with the corresponding feature, we adopt multi-channel GRU structure to avoid feature interaction at this stage. Each feature is embed with an isolated GRU, a time-series model that has a proven track record of consistent performance in EHR modeling [6]:
$\begin{array}[]{cc}\bm{h_{d_{i}}}=\mathrm{GRU}_{i}(\bm{d_{i}}),&i=1,2,\cdots,D\end{array}$ (2)
@@ 4.2.2 Dynamic Features Embedding
$\bm{h}=\mathrm{stack}(\bm{h^{\prime}_{s_{1}}},\bm{h^{\prime}_{s_{2}}},\cdots,\bm{h^{\prime}_{s_{S}}},\bm{h_{d_{1}}},\cdots,\bm{h_{d_{D}}})$ (3)
@@
4.3 Feature Confidence Learner
@@
4.3.1 Feature Missing Status Representation
@@
$\rho_{i}=\frac{\text{total observations of}\ i\text{-th feature}}{\text{total visits of all patients}}$ (4)
@@
$\tau_{i,t}=\begin{cases}0&\text{if case 1)}\\ \infty&\text{if case 2)}\\ t-t^{*}&\text{otherwise}\end{cases}$ (5)
@@
4.3.2 Missing-Aware Self-Attention
@@
${q}_{i,T}=\bm{W}_{i}^{q}\cdot{h}_{i,T}$ (6)
${k}_{i,t}=\bm{W}_{i}^{k}\cdot{h}_{i,t}$ (7)
${v}_{i,t}=\bm{W}_{i}^{v}\cdot{h}_{i,t}$ (8)
@@
$\bm{{\alpha}}_{i,*,t}={softmax}(\frac{{q}_{i,T}\bm{{k}}_{*,t}^{\top}}{\sqrt{d_{k}}})$ (9)
@@
$C_{i,t}=\begin{cases}\tanh\left(\frac{\alpha_{i,t}}{\omega_{i,t}}\right)&\text{if }\tau_{i,t}\neq\infty\\ \beta\cdot\rho_{i}&\text{if }\tau_{i,t}=\infty\end{cases}$ (10)
@@
$\omega_{i,t}={\gamma_{i}\cdot\log(e+(1-\alpha_{i,t})\cdot\tau_{i,t})}$ (11)
$\alpha_{i,t}=AvgPool(\bm{\alpha}_{i,*,t})$ (12)
@@
$\bm{\alpha^{*}}=\epsilon\cdot\bm{\alpha}+(1-\epsilon)\cdot\bm{C}$ (13)
$\bm{z}=\bm{\alpha^{*}}\bm{V}$ (14)
@@
4.4 Confidence-Aware Prototype Patient Learner
We incorporate feature confidence status into our confidence-aware module for identifying similar patient cohorts, accounting for the impact of missing features. Inspired by GRASP’s [3] graph convolutional network (GCN) framework, we further compute and integrate the similarity score within the GCN, considering the individual patient’s feature confidence to prioritize the selection of patients most similar to the subject patient.
4.4.1 Confidence-Aware Patient Similarity Measure
@@
$\phi_{i,j}=\frac{1}{(1-\zeta)\cdot\frac{1}{F}(\psi_{i,j}^{(z)}(\bm{z}_{i},\bm{z}_{j})+\zeta\cdot\psi_{i,j}^{(C)}(\bm{C}_{i},\bm{C}_{j}))}$ (15)
@@
$\psi_{i,j}^{(z)}(\bm{z}_{i},\bm{z}_{j})=||\bm{z_{i}}-\bm{z_{j}}||_{2}^{2}$ (16)
$\psi_{i,j}^{(C)}(\bm{C}_{i},\bm{C}_{j})=\sum_{k=1}^{F}\exp(1-{{C}_{i,k}})\cdot\exp(1-{{C}_{j,k}})$ (17)
4.4.2 Prototype Patients Cohort Discovery
@@
$\bm{\mathcal{G}^{*}}={\rm MLP(GCN}({\mathrm{concat}(\bm{z},\bm{\mathcal{G}})},\bm{A}))$ (18)
@@
4.4.3 Prototype Representation Fusion
@@
$\bm{z_{i}^{*}}=\eta\cdot\bm{\mathcal{G}_{i}}+(1-\eta)\cdot\bm{z_{i}}$ (19)
@@ 4.5 Prediction Layers
5 Experimental Setups
5.1 Benchmarked Real-World Datasets
We employ MIMIC-III, MIMIC-IV, PhysioNet Challenge 2012, and eICU datasets datasets for benchmarking. All 4 datasets are split into 70% training set, 10% validation set and 20% test set with stratified shuffle split strategy based on patients’ end-stage mortality outcome. By default, we use the Last Observation Carried Forward (LOCF) imputation method [23].
  1. 1.
    MIMIC-III [24] (Medical Information Mart for Intensive Care) is a large, freely-available database comprising information such as demographics, vital sign measurements made at the bedside, laboratory test results, procedures, medications, caregiver notes, imaging reports, and mortality.
  2. 2.
    MIMIC-IV [25] as MIMIC-III dataset’s subsequent iteration, stands as an evolved manifestation of the MIMIC-III database, encompassing data updates and partial table reconstructions. We extracted patient EHR data following [26].
  3. 3.
    PhysioNet Challenge 2012 [27] (Challenge-2012) covers records from 12,000 adult ICU stays. The challenge is designed to promote the development of effective algorithms for predicting in-hospital mortality based on data from the first 48 hours of ICU admission. We utilize 35 lab test features and 5 demographic features in Challenge-2012 dataset.
  4. 4.
    eICU [28] is a large-scale, multi-center ICU database derived from over 200,000 ICU admissions across the United States between 2014 and 2015. 12 lab test features and 2 demographic features are adopted.

### v6

# PRISM: Mitigating EHR Data Sparsity via Learning from Missing Feature Calibrated Prototype Patient Representations

@@ Abstract.
Electronic Health Records (EHRs) contain a wealth of patient data; however, the sparsity of EHRs data often presents significant challenges for predictive modeling. Conventional imputation methods inadequately distinguish between real and imputed data, leading to potential inaccuracies of patient representations. To address these issues, we introduce PRISM, a framework that indirectly imputes data by leveraging prototype representations of similar patients, thus ensuring compact representations that preserve patient information. PRISM also includes a feature confidence learner module, which evaluates the reliability of each feature considering missing statuses. Additionally, PRISM introduces a new patient similarity metric that accounts for feature confidence, avoiding over-reliance on imprecise imputed values. Our extensive experiments on the MIMIC-III, MIMIC-IV, PhysioNet Challenge 2012, eICU datasets demonstrate PRISM’s superior performance in predicting in-hospital mortality and 30-day readmission tasks, showcasing its effectiveness in handling EHR data sparsity. For the sake of reproducibility and further research, we have publicly released the code at https://github.com/yhzhu99/PRISM.
electronic health record; data mining; deep learning
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USA.booktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAisbn: 979-8-4007-0436-9/24/10doi: 10.1145/3627673.3679521ccs: Applied computing Health informaticsccs: Information systems Data mining
1. Introduction
@@
Electronic Health Records (EHR) have become indispensable in modern healthcare, offering a rich source of data that chronicles a patient’s medical history. Over recent years, machine learning techniques have gained significant attention for their ability to leverage time-series EHR data, which represented as temporal sequences of high-dimensional clinical variables (Choi et al., 2016), can significantly inform and enhance clinical decision-making. Such applications range from predicting the survival risk of patients (Ma et al., 2020b; Zhang et al., 2021; Liao et al., 2024) to forecasting early mortality outcomes (Yan et al., 2020; Ma et al., 2022; Gao et al., 2024; Zhu et al., 2023).
@@
Working with time-series EHR data presents challenges due to its inherent sparsity. Factors such as data corruption (Chen and Zhang, 2020), expensive examinations (Ford and Ford, 2000), and safety considerations (Zhou et al., 2020) result in missing observations; for instance, not all indicators are captured during every patient visit (Che et al., 2018). Imputed values, while necessary, are not genuine reflections of a patient’s condition and can introduce noise, diminishing model accuracy (Zhang et al., 2022). Given that most machine learning models cannot process NaN (Not a Number) inputs, this sparsity necessitates imputation, complicating EHR predictive modeling. While most existing works tackling EHR sparsity have tried to perform the imputation task directly on raw data, based on modeling the health status trajectory of the whole training set, this approach is similar to that of matrix completion methods, such as MICE (Van Buuren and Groothuis-Oudshoorn, 2011), non-negative matrix factorization (NMF) (Wang and Zhang, 2012), and compressed sensing (Lopes, 2013). However, they fail to capture the temporal interactions of longitudinal EHR data for each patient. Also, previous EHR-specific models with strategies of recalibrating patient representations based on attentive feature importance (Ma et al., 2020a, 2023) exhibit an issue of inadvertently prioritizing imputed features, potentially introducing inaccuracies (Liu et al., 2023). Thus, addressing the sparsity in time-series EHR data requires a focus on capturing relevant feature representations across visits within a patient or among patients that are essential for predictive purposes. The primary objective is to discern and highlight crucial features while diminishing the impact of irrelevant, redundant, or missing ones.
@@
Intuitively, incorporating knowledge from similar patients as an indirect method of imputation has the potential to enhance patient representations in the context of sparse EHR data. Such a knowledge-driven approach mirrors the real-world clinical reasoning processes, harnessing patterns observed in related patient cases (Suo et al., 2018). Regarding the identifying similar patient process, however, existing models also face a significant limitation: they cannot distinguish between actual and imputed data (Zhang et al., 2021, 2022). Consider two patients who have the same lab test feature value. For patient A, this value originates from the actual data, while for patient B, it is an imputed value. Current similarity metrics, whether they are L1, L2 distance, cosine similarity (Lee et al., 2015), handcrafted metrics (Huang et al., 2019), or learning-based methods (Suo et al., 2018), interpret these values identically. This results in a potentially misleading perception of similarity.
@@
Given these insights, we are confronted with a pressing challenge: How can we effectively mitigate the sparsity issue in EHR data caused by missing recorded features while ensuring a compact patient representation that preserves patient information? (Si et al., 2021)
@@
  • Methodologically, we propose PRISM, a framework for learning prototype representations of similar patients, designed to mitigate EHR data sparsity. We design the feature confidence learner that evaluates and calibrates the reliability of each feature by examining its absence and associated confidence level. We also introduce the confidence-aware prototype patient learner with enhanced patient similarity measures that differentiates between varying feature confidence levels. Compared to existing SOTA baselines, i.e. GRASP (Zhang et al., 2021) and M3Care (Zhang et al., 2022), PRISM provides a refined feature calibration, and further missing-aware similarity measure helps to identify more related patient representation, with missing feature status elaborately taking into account.
@@ 2. Related Work
2.1. Direct Imputation
Direct imputation methods aim to estimate missing features or incorporate missing information directly. Traditional matrix imputation techniques, such as MICE (Van Buuren and Groothuis-Oudshoorn, 2011), non-negative matrix factorization (NMF) (Wang and Zhang, 2012), compressed sensing (Lopes, 2013), or naive zero, mean, or median imputation, rely on similar rows or columns to fill in missing data. However, these methods often operate under the assumption that patient visits are independent and features are missing at random. GRU-D (Che et al., 2018) adopts a more targeted approach by introducing missing statuses into the GRU network. By utilizing time interval and missing mask information, GRU-D treats missing data as “Informative Missing”. Extending these capabilities, ConCare (Ma et al., 2020b) and AICare (Ma et al., 2023) first apply directly imputed EHR data, then incorporate multi-head self-attention mechanisms to refine feature embeddings. This ensures contextual relevance across diverse healthcare situations, regardless of data completeness.
2.2. Indirect Imputation
As demonstrated in GRASP (Zhang et al., 2021) and M3Care (Zhang et al., 2022), they emphasize the use of similar patient representations to derive meaningful information with the insight that the information observed from similar patients can be utilized as guidance for the current patient’s prognosis (Zhang et al., 2021). However, accurately measuring patient similarity is intrinsically challenging, especially when features might be imputed with potentially misleading information. Many traditional works, such as  (Lee et al., 2015) and  (Huang et al., 2019), have resorted to fixed formulas like cosine similarity and Euclidean distance to gauge patient similarity. While these methods are straightforward, they often suffer from scalability and performance limitations. A more dynamic approach is seen in  (Suo et al., 2018), which adopts metric learning with triplet loss. This technique focuses on learning the relative distances between patients, where distances have an inverse correlation with similarity scores.
@@ 2.2. Indirect Imputation
3. Problem Formulation
3.1. EHR Datasets Formulation
@@
3.2. Predictive Objective Formulation
@@ 3.3. Notation Table
Table 1. Notations symbols and their descriptions
@@
4. Methodology
4.1. Overview
Figure 1. Overall model architecture of our proposed method PRISM. “Rep.” means “Representation”.
@@ 4.1. Overview
4.2. Feature-Isolated Embedding Module
@@
4.2.1. Static Features Embedding
@@
(1) $\begin{array}[]{cc}\bm{h_{s_{i}}}=\mathrm{MLP}_{i}(\bm{s_{i}}),&i=1,2,\cdots,S\end{array}$
@@
4.2.2. Dynamic Features Embedding
To ensure that each feature’s individual statistics, e.g. missing status can be incorporated with the corresponding feature, we adopt multi-channel GRU structure to avoid feature interaction at this stage. Each feature is embed with an isolated GRU, a time-series model that has a proven track record of consistent performance in EHR modeling (Gao et al., 2024):
(2) $\begin{array}[]{cc}\bm{h_{d_{i}}}=\mathrm{GRU}_{i}(\bm{d_{i}}),&i=1,2,\cdots,D\end{array}$
@@ 4.2.2. Dynamic Features Embedding
(3) $\bm{h}=\mathrm{stack}(\bm{h^{\prime}_{s_{1}}},\bm{h^{\prime}_{s_{2}}},\cdots,\bm{h^{\prime}_{s_{S}}},\bm{h_{d_{1}}},\cdots,\bm{h_{d_{D}}})$
@@
4.3. Feature Confidence Learner
@@
4.3.1. Feature Missing Status Representation
@@
(4) $\rho_{i}=\frac{\text{total observations of}\ i\text{-th feature}}{\text{total visits of all patients}}$
@@
(5) $\tau_{i,t}=\begin{cases}0&\text{if case 1)}\\ \infty&\text{if case 2)}\\ t-t^{*}&\text{otherwise}\end{cases}$
@@
4.3.2. Missing-Aware Self-Attention
@@
(6) ${q}_{i,T}=\bm{W}_{i}^{q}\cdot{h}_{i,T}$
(7) ${k}_{i,t}=\bm{W}_{i}^{k}\cdot{h}_{i,t}$
(8) ${v}_{i,t}=\bm{W}_{i}^{v}\cdot{h}_{i,t}$
@@
(9) $\bm{{\alpha}}_{i,*,t}={softmax}(\frac{{q}_{i,T}\bm{{k}}_{*,t}^{\top}}{\sqrt{d_{k}}})$
@@
(10) $C_{i,t}=\begin{cases}\tanh\left(\frac{\alpha_{i,t}}{\omega_{i,t}}\right)&\text{if }\tau_{i,t}\neq\infty\\ \beta\cdot\rho_{i}&\text{if }\tau_{i,t}=\infty\end{cases}$
@@
(11) $\omega_{i,t}={\gamma_{i}\cdot\log(e+(1-\alpha_{i,t})\cdot\tau_{i,t})}$
(12) $\alpha_{i,t}=AvgPool(\bm{\alpha}_{i,*,t})$
@@
(13) $\bm{\alpha^{*}}=\epsilon\cdot\bm{\alpha}+(1-\epsilon)\cdot\bm{C}$
(14) $\bm{z}=\bm{\alpha^{*}}\bm{V}$
@@
4.4. Confidence-Aware Prototype Patient Learner
We incorporate feature confidence status into our confidence-aware module for identifying similar patient cohorts, accounting for the impact of missing features. Inspired by GRASP’s (Zhang et al., 2021) graph convolutional network (GCN) framework, we further compute and integrate the similarity score within the GCN, considering the individual patient’s feature confidence to prioritize the selection of patients most similar to the subject patient.
4.4.1. Confidence-Aware Patient Similarity Measure
@@
(15) $\phi_{i,j}=\frac{1}{(1-\zeta)\cdot\psi_{i,j}^{(z)}(\bm{z}_{i},\bm{z}_{j})+\zeta\cdot\psi_{i,j}^{(C)}(\bm{C}_{i},\bm{C}_{j})}$
@@
(16) $\psi_{i,j}^{(z)}(\bm{z}_{i},\bm{z}_{j})=\frac{1}{F}\|\bm{z}_{i}-\bm{z}_{j}\|_{2}^{2}$
(17) $\psi_{i,j}^{(C)}(\bm{C}_{i},\bm{C}_{j})=\frac{1}{F}\sum_{k=1}^{F}\exp(1-C_{i,k})\cdot\exp(1-C_{j,k})$
4.4.2. Prototype Patients Cohort Discovery
@@
(18) $\bm{\mathcal{G}^{*}}={\rm MLP(GCN}({\mathrm{concat}(\bm{z},\bm{\mathcal{G}})},\bm{A}))$
@@
4.4.3. Prototype Representation Fusion
@@
(19) $\bm{z_{i}^{*}}=\eta\cdot\bm{\mathcal{G}_{i}}+(1-\eta)\cdot\bm{z_{i}}$
@@ 4.5. Prediction Layers
5. Experimental Setups
5.1. Benchmarked Real-World Datasets
We employ MIMIC-III, MIMIC-IV, PhysioNet Challenge 2012, and eICU datasets for benchmarking. All 4 datasets are split into 70% training set, 10% validation set and 20% test set with stratified shuffle split strategy based on patients’ end-stage mortality outcome. By default, we use the Last Observation Carried Forward (LOCF) imputation method (Woolley et al., 2009).
  1. (1)
    MIMIC-III (Johnson et al., 2016) (Medical Information Mart for Intensive Care) is a large, freely-available database comprising information such as demographics, vital sign measurements made at the bedside, laboratory test results, procedures, medications, caregiver notes, imaging reports, and mortality.
  2. (2)
    MIMIC-IV (Johnson et al., 2023) as MIMIC-III dataset’s subsequent iteration, stands as an evolved manifestation of the MIMIC-III database, encompassing data updates and partial table reconstructions. We extracted patient EHR data following (Harutyunyan et al., 2019).
  3. (3)
    PhysioNet Challenge 2012 (Silva et al., 2012) (Challenge-2012) covers records from 12,000 adult ICU stays. The challenge is designed to promote the development of effective algorithms for predicting in-hospital mortality based on data from the first 48 hours of ICU admission. We utilize 35 lab test features and 5 demographic features in Challenge-2012 dataset.
  4. (4)
    eICU (Pollard et al., 2018) is a large-scale, multi-center ICU database derived from over 200,000 ICU admissions across the United States between 2014 and 2015. 12 lab test features and 2 demographic features are adopted.